CAT Chat Minutes

July 14, 2000

Information and Follow-up Items

1. When will the EPICS control of the liquid nitrogen valve be made available to the beamline?

The EPICS controls have been tested. We expect to have the controls available to the CATs by August.

2. With regards to the fume hoods, has anyone considered different filters?

There is a study being done right now by engineering and B. Glagola to determine if the hoods can be upgraded to accommodate the needs of the user community.

 

General Information

The switchgear maintenance will take place today, Friday, July 14th. This will begin at 5:00 pm until 6:30 pm and will affect Sectors 1, 2 and 3, and LOM 431. The floor coordinators have contacted SRI-CAT of this particular maintenance.

The low temperature hot water system will be shutdown on Tuesday, July 18th for maintenance. Many leaks have been occurred in the past; therefore, we will be testing new gaskets to prevent leaks in the future.

Mohan welcomed back Yeldez Amer from her long recovery. For the immediate future, Yeldez will concentrate in the QA area and work with documentation required by the APS Operations Division.

APS Status Information handouts were provided to the group. This information will discuss how to retrieve some of the EPICS Process Variables on the status of the Storage Ring. This web page is updated periodically and can be found at:

http://www.aps.anl.gov/xfd/operations/APS_Status_Information3.html

If you're interested in additional process variables, which are not listed in these documents, contact your floor coordinator and he/she will provide you with one.

The LOM chemistry lab fume hoods are in the process of being reevaluated. A few areas to be looked into are measured performance values, hood controls, noise control, high flow rates, protection and certification of the hoods. More data will be ready within two weeks.